The end zone of a football field is 10 yards deep and 53 yards across. Between the two end zones is the playing field, which is 100 yards long and 53 yards wide. What is the length of the diagonal from the back of one end zone to the back of the other?
Answer by nerdybill(7384)

The end zone of a football field is 10 yards deep and 53 yards across. Between the two end zones is the playing field, which is 100 yards long and 53 yards wide. What is the length of the diagonal from the back of one end zone to the back of the other?
.
Let d = length of diagonal
then
d^2 = 53^2 + (10+100+10)^2
d^2 = 53^2 + 120^2
d^2 = 2809 + 14400
d^2 = 17209
d = 131.18 yards